The amount of Zinc required to produce 1.12 ml of H2 at STP on treatment with dilute HCL will be? plz ansr soon Thanks in advance

Grade:11

3 Answers

Beenu Mathew
7 years ago
HI,
Zn + 2HCl -----> H2 + ZnCl2
Now it is given in the question that 1.12 ml at STP is formed
so????
look at the equation carefully
you will see 
1 mole of Zn gives you 1 mole of H2
so ...yea its on this basis we will be solving the rest of this question
1.12 ml of H2 is how many moles
1.12/22.4 = .05 moles
so 0.05 moles of Zn is required
0.05 * 65.4 = 3.27g (by the way 65.4g is the molar mass of Zn.,,,,checked it on google ;) )
THANKS!!
BEST OF LUCK 
CHEERS!!
 
Akash Biswas
6 years ago
Zn + 2HCl --» ZnCl2 + H21 Mole of zinc reacts to produce 1 mole of hydrogen gas ..So 1.12 ml of hydrogen gas contains 1.12 ml/22400 ml = 0.5* 10 ^ -4Thus 0.5* 10 ^ -4 mole if hydrogen gas is obtained from 0.5* 10 ^ -4 mole of Zinc So amount of zinc required to produce 1.12 ml of hydrogen gas is Molecular weight of Zn*0.5* 10 ^ -4=65.38*0.5* 10 ^ -4= 32.5 * 10 ^ -4 g
Yashvi
4 years ago
Zn + 2HCl= ZnCl+ H2(g)
1 mol(65 g - atomic weight) Zn required for 1 mol H2(22.4l=22400ml)
If 65g Zn - 22400ml
Then x g Zn - 1.12 ml
1.12×65/22400= 0.00325 i.e. 3.25×10-4